Grade 12Organic Chemistry

tuning forks when sounded tigether produce 3 beats per sceonds. on loading one of these forks with little wax and sounding them together , 20 beats are heard in 4 seconds. calculate the frequency of the loaded tuning fork if the frequency of the other fork is 368 HZ. . explain your answer.

sheikh
7 Years agoGrade 12

1 Answer

V Prakash
7 Years ago
when fork is waxed freq. must decrease. fbeats=fa-fb.  fb=368. when fa decreases fbeats must decrease but its increasing. hence our equation must be fbeats=fb-fa.  3=fa-368. fa=371. when loaded beats per second are 20/4=5. then 5=fa(loaded)-368.  hence freq. of loaded fork is 373